the official height-to-width ratio of the united states flag is 1 : 1.9 . if a united states flag is 9.5 feet wide, how high should it be?

Answer by jlshires(17)

If the ratio of the flag is 1:1.9 h:w, and the flag is 9.5 w, you have to turn it into a comparable fraction.
h/w = 1/1.9 = x/9.5 you then cross multiply so:
(1)(9.5) = x 1.9 which is equal to
9.5 = 1.9x divide both sides by 1.9
5 = x
The height should be 5 feet
(** Always make sure you include the unit of measurement in your answer)